# Potential Fields as Scene Affordance for Behavior Change-Based Visual Risk Object Identification

### v1

@@ Abstract
We study behavior change-based visual risk object identification (Visual-ROI), a critical framework designed to detect potential hazards for intelligent driving systems. Existing methods often show significant limitations in spatial accuracy and temporal consistency, stemming from an incomplete understanding of scene affordance. For example, these methods frequently misidentify vehicles that do not impact the ego vehicle as risk objects. Furthermore, existing behavior change-based methods are inefficient because they implement causal inference in the perspective image space. We propose a new framework with a Bird’s Eye View (BEV) representation to overcome the above challenges. Specifically, we utilize potential fields as scene affordance, involving repulsive forces derived from road infrastructure and traffic participants, along with attractive forces sourced from target destinations. In this work, we compute potential fields by assigning different energy levels according to the semantic labels obtained from BEV semantic segmentation. We conduct thorough experiments and ablation studies, comparing the proposed method with various state-of-the-art algorithms on both synthetic and real-world datasets. Our results show a notable increase in spatial and temporal consistency, with enhancements of 20.3% and 11.6% on the RiskBench dataset, respectively. Additionally, we can improve computational efficiency by 88%. We achieve improvements of 5.4% in spatial accuracy and 7.2% in temporal consistency on the nuScenes dataset. For more qualitative results, please visit our project webpage: project webpage.
@@ I Introduction
Intelligent driving systems such as advanced driver assistance systems (ADAS) attract significant attention in academia and industry, aiming to provide immediate alerts to help reduce the number of road accidents. One of the indispensable technologies is visual risk object identification (Visual-ROI), which involves localizing potential hazards and estimating the corresponding risk scores. The community has explored a wide range of approaches, including collision prediction [1, 2, 3, 4, 5], trajectory prediction and collision checking [6, 7, 8], object importance estimation [9, 10, 11, 12, 13], human gaze prediction [14, 15, 16, 17, 18], and behavior change-based prediction [19, 20, 21, 22, 23].
@@ II-A Visual Risk Object Identification
Visual risk object identification (Visual-ROI) [9, 14, 1, 10, 11, 2, 15, 29, 3, 12, 4, 6, 20, 19, 16, 13, 5, 7, 8, 17, 18, 22, 30, 23] is an essential technology for the development of intelligent driving systems. The field can be categorized into the following four paradigms. First, predicting objects that potentially involve in collision is defined as risk object [1, 2, 3, 4, 5, 6, 7, 8]. Second, risk objects are defined based on human annotators’ subjective assessment [9, 10, 11, 12, 13, 31]. Third, objects focused by human gaze are defined as risk objects [14, 15, 16, 17, 18]. Fourth, objects influencing either the driver’s or ego vehicle’s are defined as risk objects [20, 24, 22, 23, 19].
@@ III-D Behavior Change-based Visual Risk Object Identification
In this work, we make the following two changes to [20], referred to as PF+BCP. First, we conduct causal inference in the bird’s-eye view instead of the perspective view. Precisely, given a rendered potential field and an object tracklet, we remove the corresponding repulsive force to simulate a potential field without the tracklet. Second, we train a behavior change predictor using the rendered potential fields as input.
@@ IV-B Baselines
TP+BCP: We incorporate a target point predictor into BCP and use the same mechanism to define their risk scores.
@@ IV-C Metrics
We evaluate the performance of Visual-ROI models with two types of metrics: Spatial Accuracy and Temporal Consistency.
@@ V-A Visual Risk Object Identification
We conduct extensive experiments on the RiskBench dataset. We show the proposed method PF+BCP achieves state-of-the-art performance compared to existing Visual-ROI methods, as shown in Table V-A. Our method has shown remarkable effectiveness, yielding a 20.3% improvement in the F1 score over the BCP [20] on the RiskBench dataset. Additionally, we achieve a 2.6% improvement in the F1 score over BS+BCP, highlighting the benefit of incorporating attractive and repulsive forces into the model. Note that the improvements in both recall and precision further validate the robustness of our approach.
In terms of temporal consistency, PF+BCP shows an improvement in wMOTA by 11.6% compared to BCP and 2.3% compared to BS+BCP. The improvements in PIC further validate that the proposed method enhances spatial accuracy and temporal consistency over existing models. A similar trend is observed in the OT-F1-T results as well. As shown in Table V-A, PF+BCP improves the OT-F1-T score by 13–15% compared to BCP, demonstrating that the proposed method provides stable predictions near the critical frames.
@@ V-B Scenario-based Analysis
We perform a scenario-based analysis to assess model performance across various traffic conditions. As shown in Table V-B, PF+BCP demonstrates fewer false positives compared to all baselines in the opposite-lane scenario. This strong performance is due to its focus on objects that may influence the ego vehicle’s progress toward its target (attractive force). Moreover, the modeling of repulsive force helps identify non-afforded actions toward objects in the opposite lane.
PF+BCP significantly outperforms other methods in scenarios involving the various actors, as shown in Table V-B. The results from the Closest Actors demonstrate that the PF+BCP method effectively identifies risk traffic participants in the vicinity of the ego vehicle. The Truck scenario demonstrates superior performance across all methods, likely because BEV semantic segmentation can detect large objects. In contrast, the Pedestrians scenarios show less satisfactory results, as BEV-SEG struggles to accurately capture pedestrian features. This limitation leads to errors in the causal inference stage and impacts the performance of Visual-ROI. We find that small objects, like pedestrians, pose a unique challenge, highlighting the need for further advancement in BEV semantic segmentation. Our scenario-based analysis demonstrates the versatility and effectiveness of our approach across various driving conditions, while also identifying the need for improvement in small object detection.
@@ V-C Ablation Studies
We conduct extensive ablation studies to justify our design choices built upon BCP (ID 1) [20]. The results are presented in Table V-B. Adding a target point predictor to BCP (ID 2) improves OT-F1 by 8%, highlighting the importance of modeling ego vehicle’s intention. Incorporating BEV semantic segmentation (ID 3) further enhances OT-F1 by nearly 10%, demonstrating that rich semantic context aids in risk object identification. In ID 4, we assess the importance of the repulsive force without a target point predictor and find its performance comparable to ID 3. We conjecture that the lack of attractive force may not be fully aware of the risk object toward the target location, increasing the number of misidentifications. Finally, combining all the designed modules (ID 5) yields the best results, i.e., an 20% improvement over BCP.
@@ V-D Real-world Evaluation
We conduct a real-world evaluation on the nuScenes [28] dataset. As shown in Table V-D, PF+BCP outperforms the baselines BP, BCP and BS+BCP in both spatial accuracy and temporal consistency metrics. Specifically, PF+BCP enhances OT-F1 by almost 30% and 5.4% over BP and BCP, respectively. The proposed method also decreases PIC by 10.1% and 6.3% compared to BP and BCP. Moreover, with our proposed method, we improve wMOTA by 17.9% and 7.2% over BP and BCP, respectively.
@@
Fig. 3 presents the qualitative results of our model across various traffic scenarios. We observe that BCP exhibits several false alarms, while PF+BCP reliably identifies risk objects, such as crossing pedestrians. Additionally, the proposed method avoids misidentifying vehicles in the opposite lane as risk objects. These results confirm that our method effectively identifies potential risks and significantly reduces false alarms in complex scenarios, ensuring high reliability in real-world applications. Additional qualitative results are available in the supplementary video and on our project page.
@@ VI Conclusion
In this work, we propose using potential fields as a new representation for scene affordance.
We demonstrate that integrating potential fields can address the limitations of existing behavior change-based Visual-ROI algorithms, specifically, spatial inaccuracy, temporal inconsistency and computational inefficiency.
Through comprehensive experiments and ablation studies on the RiskBench dataset, we achieve relative enhancements of 11.6% in wMOTA and 20.3% in the F1 score compared to strong baselines. Moreover, the proposed method can speed up causal inference by 88%.
We also evaluate our approach on the nuScenes dataset, achieving a 5.4% improvement in wMOTA and a 7.2% improvement in the F1 score.
This demonstrates that potential fields as scene affordance can facilitate behavior change-based Visual-ROI in real-world scenarios.
These results underscore the vital role of scene affordance in advancing the safety and efficiency of intelligent driving systems.
Limitation and Future Work:
The effectiveness of the potential fields in our approach is highly dependent on the quality of BEV semantic segmentation.
For instance, Fig. 4 shows failure cases due to imperfect BEV segmentation.
Manually defining repulsive and attractive force constants limits the broader applicability of our method.
To address this, we plan to explore alternative potential field rendering strategies, such as the approach in [77], and investigate more diverse scene affordances, like drivable areas and lane concepts, for integration into decision-making tasks.

### v2

@@ Abstract
We study behavior change-based visual risk object identification (Visual-ROI), a critical framework designed to detect potential hazards for intelligent driving systems. Existing methods often show significant limitations in spatial accuracy and temporal consistency, stemming from an incomplete understanding of scene affordance. For example, these methods frequently misidentify vehicles that do not impact the ego vehicle as risk objects. Furthermore, existing behavior change-based methods are inefficient because they implement causal inference in the perspective image space. We propose a new framework with a Bird’s Eye View (BEV) representation to overcome the above challenges. Specifically, we utilize potential fields as scene affordance, involving repulsive forces derived from road infrastructure and traffic participants, along with attractive forces sourced from target destinations. In this work, we compute potential fields by assigning different energy levels according to the semantic labels obtained from BEV semantic segmentation. We conduct thorough experiments and ablation studies, comparing the proposed method with various state-of-the-art algorithms on both synthetic and real-world datasets. Our results show a notable increase in spatial accuracy and temporal consistency, with enhancements of 20.3% and 11.6% on the RiskBench dataset, respectively. Additionally, we can improve computational efficiency by 88%. We achieve improvements of 5.4% in spatial accuracy and 7.2% in temporal consistency on the nuScenes dataset. For more qualitative results, please visit our project webpage: project webpage.
@@ I Introduction
Intelligent driving systems such as advanced driver assistance systems attract significant attention in academia and industry, aiming to provide immediate alerts to reduce the number of road accidents. One of the indispensable technologies is visual risk object identification (Visual-ROI), which involves localizing potential hazards and estimating the corresponding risk scores. The community has explored a wide range of approaches, including collision prediction [1, 2, 3, 4, 5], trajectory prediction and collision checking [6, 7, 8], object importance estimation [9, 10, 11, 12, 13], human gaze prediction [14, 15, 16, 17, 18], and behavior change-based prediction [19, 20, 21, 22, 23].
@@ II-A Visual Risk Object Identification
Visual risk object identification (Visual-ROI) [9, 14, 1, 10, 11, 2, 15, 29, 3, 12, 4, 6, 20, 19, 16, 13, 5, 7, 8, 17, 18, 21, 30, 22] is an essential technology for the development of intelligent driving systems. The field can be categorized into the following four paradigms. First, predicting objects that potentially involve in collision is defined as risk object [1, 2, 3, 4, 5, 6, 7, 8]. Second, risk objects are defined based on human annotators’ subjective assessment [9, 10, 11, 12, 13, 31]. Third, objects focused by human gaze are defined as risk objects [14, 15, 16, 17, 18]. Forth, objects influencing either the driver’s or ego vehicle’s behavior are defined as risk objects [20, 24, 21, 22, 19].
@@ III-D Behavior Change-based Visual Risk Object Identification
In this work, we make the following two changes to [20], referred to as PF+BCP. First, we conduct causal inference in the BEV instead of the perspective view. Precisely, given a rendered potential field and an object tracklet, we remove the corresponding repulsive force to simulate a potential field without the tracklet. Second, we train a behavior change predictor using the rendered potential fields as input.
@@ IV-B Baselines
TP+BCP: We integrate from the target point predictor into BCP, using the same mechanism to define risk scores. Here, is represented as a 2-dimensional vector as input.
@@ IV-C Metrics
We evaluate Visual-ROI models with two types of metrics: Spatial Accuracy and Temporal Consistency.
@@ V-A Visual Risk Object Identification
We conduct extensive experiments on the RiskBench dataset [25]. We show the proposed method PF+BCP achieves state-of-the-art performance compared to existing Visual-ROI methods, as shown in Table V-A. Our method has shown remarkable effectiveness, yielding a 20.3% improvement in the F1 score over the BCP [20] on the RiskBench dataset. Additionally, we achieve a 2.6% improvement in the F1 score over BS+BCP, highlighting the benefit of incorporating attractive and repulsive forces into the model. Note that the improvements in both recall and precision further validate the robustness of our approach.
In terms of temporal consistency, PF+BCP shows an improvement in wMOTA by 11.6% compared to BCP and 2.3% compared to BS+BCP. The improvements in PIC further validate that the proposed method enhances spatial accuracy and temporal consistency over existing models. A similar trend is also observed in the OT-F1-T results. As shown in Table V-A, PF+BCP improves the OT-F1-T score by 13–15% compared to BCP, demonstrating that the proposed method provides stable predictions near the critical frames.
@@ V-B Scenario-based Analysis
We perform a scenario-based analysis to assess model performance in various traffic conditions. As shown in Table V-B, PF+BCP demonstrates fewer false positives compared to all baselines in the opposite-lane scenario. This strong performance is due to its focus on objects that may influence the ego vehicle’s progress toward its target (attractive force). Moreover, the modeling of repulsive force helps identify non-afforded actions toward objects in the opposite lane.
PF+BCP significantly outperforms other methods in scenarios involving the various actors, as shown in Table V-B. The results from the Closest Actors demonstrate that the PF+BCP method effectively identifies risk traffic participants in the vicinity of the ego vehicle. The Truck scenario demonstrates superior performance across all methods, likely because BEV semantic segmentation can detect large objects. In contrast, the Pedestrians scenarios show less satisfactory results, as BEV-SEG struggles to accurately capture pedestrian features. This limitation leads to errors in the causal inference stage and impacts the performance of Visual-ROI. We find that small objects, such as pedestrians, pose a unique challenge, highlighting the need for further advancement in BEV semantic segmentation. Our scenario-based analysis demonstrates the versatility and effectiveness of our approach across various driving conditions, while also identifying the need for improvement in small object detection.
@@ V-C Ablation Studies
We conduct extensive ablation studies to justify our design choices built upon BCP (ID 1) [20]. The results are presented in Table V-B. Adding a target point predictor to BCP (ID 2) improves OT-F1 by 8%, highlighting the importance of modeling ego vehicle’s intention. Incorporating BEV semantic segmentation (ID 3) further enhances OT-F1 by nearly 10%, demonstrating that rich semantic context aids in risk object identification. In ID 4, we assess the importance of the repulsive force without attractive force and find its performance comparable to ID 3. We conjecture that the lack of attractive force may not be fully aware of the risk object toward the target location, increasing the number of misidentifications. Finally, combining all the designed modules (ID 5) yields the best results, i.e., an 20% improvement over BCP.
@@ V-D Real-world Evaluation
We conduct a real-world evaluation on the nuScenes [28] dataset. As shown in Table V-D, PF+BCP outperforms the baselines BP, BCP and BS+BCP in both spatial accuracy and temporal consistency metrics. Specifically, compared to BP and BCP, PF+BCP improves OT-F1 by 29% and 5.4%, reduces PIC by 10.1% and 6.3%, and enhances wMOTA by 17.9% and 7.2%, respectively.
@@
Fig. 3 presents the qualitative results across various traffic scenarios. We observe that BCP exhibits several false alarms, while PF+BCP reliably identifies risk objects, such as crossing pedestrians. Obviously, the proposed method avoids misidentifying vehicles in the opposite lane as risk objects. These results confirm that our method effectively identifies potential risks and significantly reduces false alarms in complex scenarios, ensuring reliability in real-world applications.
@@ VI Conclusion
In this work, we propose using potential fields as a new representation for scene affordance. We demonstrate that integrating potential fields can address the limitations of existing behavior change-based Visual-ROI algorithms, specifically, spatial inaccuracy, temporal inconsistency and computational inefficiency. Through comprehensive experiments and ablation studies on the RiskBench dataset, we achieve relative enhancements of 11.6% in wMOTA and 20.3% in the F1 score compared to strong baselines. Moreover, the proposed method can speed up causal inference by 88%. We also evaluate our approach on the nuScenes dataset, achieving a 5.4% improvement in wMOTA and a 7.2% improvement in the F1 score. These results underscore the vital role of scene affordance in advancing the safety and efficiency of intelligent driving systems.
Limitation and Future Work: The effectiveness of the potential fields in our approach is highly dependent on the quality of BEV semantic segmentation. For instance, Fig. 4 shows failure cases due to imperfect BEV segmentation. Manually defining repulsive and attractive force constants limits the broader applicability of our method. To address this, we plan to explore alternative potential field rendering strategies, such as the approach in [77], and investigate more diverse scene affordances, like drivable areas and lane concepts, for integration into decision-making tasks.
Acknowledgement: The work is sponsored in part by the Higher Education Sprout Project of the National Yang Ming Chiao Tung University and Ministry of Education (MOE), the Yushan Fellow Program Administrative Support Grant, and the National Science and Technology Council (NSTC) under grants 112-2634-F-002-006, 113-2634-F-002-007, 113-2628-E-A49-022, and Institute for Information Industry (III).